1. The hydration of carbon dioxide CO2 in red blood cells takes place due to an enzyme catalytic process, where bicarbonate ions are formed:

CO2(g) + H2O(l) ® HCO3¯ (aq) + H+(aq)

At the enzyme concentration of [E]0 = 2.3´10-6 mol/L, the following kinetic data for initial reaction rate vs. initial concentration of carbon dioxide were obtained:

[CO2], mol/L

1.25

2.50

5.00

20.00

Rate, (mol/L)s-1

2.78´10-2

5.01´10-2

8.32´10-2

1.67´10-1

Determine the Michaelis constant and the turnover number of the catalytic process.

2. The enzyme a-chymotrypsin is secreted in the pancreas of mammals and cleaves peptide bonds made between certain amino acids. Several solutions containing the small peptide N-glutaryl-1-phenylalanine-p-nitroanilide (a substrate) at different concentrations were prepared and the same amount of a-chymotrypsin was added to each one. The following data were obtained on the initial rates of the formation of product:

[S]0, mol/L

0.334

0.450

0.667

1.00

1.33

1.67

Rate, (mol/L)s-1

0.152

0.201

0.269

0.417

0.505

0.667

Determine the maximum reaction rate and the Michaelis constant for the enzyme reaction.
Answer #1

1.

The plot of 1/ rate vs 1/ [CO2] is as shown

Rate CO21/[rate] 1/[CO2] 0.0278 0.0501 0.0832 0.167 40 35.971223 19.9600798 0.4 12.0192308 0.2 5.988023950.05 0.8 35 2.5 30 y

Y intercept = \dfrac {1}{V_{max}}= 4

Hence, the maximum reaction rate

V_{max}= \dfrac {1}{4}= 0.25 \; (mol/L)s^{-1}

Slope = \dfrac {K_M}{V_{max}}

39.96= \dfrac {K_M}{0.25 \: (mol/L)s^{-1}}

Hence, the Michaelis constant

K_M = 9.99 \: (mol/L)

K_{cat} = \dfrac {V_{max}}{[E]_0}= \dfrac {0.25 \: (mol/L)s^{-1}}{2.3 \times 10^{-6 } \: mol/L}

K_{cat} = 108696 \: /s